OpenStudy (anonymous):

Express each product in simplest form 3wx/6x • 3wx/9w

OpenStudy (unklerhaukus):

start by cancelling some common factors

OpenStudy (unklerhaukus):

\[\frac{3wx}{6x} \times \frac{3wx}{9w}\] \[\frac{3w\cancel{x}}{6\cancel{x}} \times \frac{3\cancel{w}x}{9\cancel{w}}\]